Determine the equation of the line that passes thru P1(4,-8) and P2(0, ¾).
1
Expert's answer
2020-09-24T12:40:31-0400

Step 1.

Find the gradient of the line using the points given.

gradient (m)= change in y-axis / change in x- axis

=(-8- 3/4)/(4-0)

=-35/16

Step2.

Using the gradient (m), one of the given points(am using P2(0,3/4) and an unknown P3(x,y)

find the equation.

-35/16= (y-3/4)/(x-0)

=(y-3/4)/x

-35x/16= (y-3/4)

y= -35x/16 +3/4